Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y).
x + y = 15
4x + 3y = 38

Answer by Alan3354(69443)

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y).
x + y = 15
4x + 3y = 38
-------------
x + y = 15 --> y = 15-x
Sub for y in the 2nd eqn and find x.